Problem 4 Let S and T have the joint probability density function fs,T(8,t) = , 0<x<1, 52 <t<8 (a) Find marginal pdfs fs(s) acan someone help me solve both a and b? Thanks.
Answer #1

lo ele ay f6)-5, 6, 7(St)dt = 3 + dt = lns - Auga ef us, ousar 5, 20)* 89,1kst) ds , $kt > Save *4d5 :00 sp (es- 4- $ +5p6$$=) *635-**** ******* f-c0)2 4 how 4.68s: ѕү ү т ggo 1>7>0